OpenStudy (anonymous):

what is the slope intercept form or 3y+2x=6

OpenStudy (anonymous):

rearrange to get in the form y=mx+b slope is m y-intercept is b

OpenStudy (anonymous):

no i just want the answer im trying to prove a point tp someone

OpenStudy (anonymous):

Just solve for y. 3y=-2x+6 y=-2/3x+2

OpenStudy (anonymous):

thnx!